The revenue function​ R(x) and the cost function​ C(x) for a particular product are given. These functions are valid only for the specified range of values. Find the number of units that must be produced to break even.
R(x)=200x-2x^2; C(x)=-x^2+25x+4900, 0< or equal x The manufacturer must produce _______ units to break even.

Answer by Boreal(15235)

R(x)=200x-2x^2; C(x)=-x^2+25x+4900
Set them equal to each other to break even.
200x-2x^2=-x^2+25x+4900
x^2-175x+4900=0
(x-35)(x-140)=0
x=35 or 140, but x cannot be greater than 100.
check 35
7000-2450=4550.
-1225+875+4900=4550
35 units ANSWER
-----------------